OpenStudy (anonymous):

Can´t see what´s wrong? . We start with: \(\Large x2 + 6x + 5 = x2 - 1\) Factor both sides of the equation: \(\Large (x + 5)(x + 1)=(x - 1)(x + 1)\) Divide both sides by \(\Large (x + 1)\) \(\Large (x + 5)=(x - 1)\) Subtract \(\Large x\) from both sides \(\Large 5=-1\)

hartnn (hartnn):

\(x^2+6x+5=x^2−1\) gives you 6x +5 = -1 6x = -6 x= -1 x+1 = 0 since x+1 =0, you CANNOT divide both sides by x+1

hartnn (hartnn):

dividing both sides by 0, is prohibited in mathematics

hartnn (hartnn):

if you divide both sides by 0, you might end up in math jail for breaking the math rules :P